OpenStudy (anonymous):

What is the critical angle theta c for total internal reflection if the material is clear acrylic, with a refractive index of 1.5? 62 degrees, 90 degrees , 42 degrees , or 49 degrees :/

OpenStudy (michele_laino):

here we have to apply the law of Snell, so we can write this: \[n\sin {\theta _c} = 1\]

OpenStudy (michele_laino):

then: \[\large \sin {\theta _c} = \frac{1}{n} = \frac{1}{{1.5}}\]

OpenStudy (anonymous):

so we get 0.666666667?

OpenStudy (michele_laino):

ok! Now we have to compute this: \[{\theta _c} = \arcsin \left( {0.667} \right) = ...\]

OpenStudy (michele_laino):

\[\large {\theta _c} = \arcsin \left( {0.667} \right) = ...\]

OpenStudy (anonymous):

41.835? so our solution is 42 degrees?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay! thank you:)

OpenStudy (michele_laino):

:)